OpenStudy (anonymous):

Myininaya, Satellite - Please help me solve these two Trigonometric Identities! Thanks! cos4(x) = cos^4(x) - 6sin^2(x) cos^2(x) + sin^4(x) sec^2(x) = 2sec(x) / sec(x) + 1

OpenStudy (anonymous):

i am confused because it looks like you have \[\cos^4(x)\]on both sides

OpenStudy (anonymous):

the second cos^4x, the 4 is a power

OpenStudy (anonymous):

like, cos to the fourth of x

OpenStudy (anonymous):

ooh and is the first one \[\cos(4x)\]?

OpenStudy (anonymous):

Yes!! You got it.

OpenStudy (anonymous):

I'm sorry

OpenStudy (anonymous):

ok then i think it will work out fine if we use the "double angle formula" twice. \[\cos(2x)=\sin^2(x)-\cos^2(x)\] among other things, so maybe if we try \[\cos(4x)=\sin^2(2x)-\cos^2(2x)\] as a start and then use it again we will get what we want. let me try to write it on paper first before i say something incorrect

OpenStudy (anonymous):

Yes, good idea. I see exactly how you figured that out. However, I am completely stuck after that.

OpenStudy (anonymous):

SO we are working with the left side ?

OpenStudy (anonymous):

yes this is what i have on paper so far \[\cos(4x)=\cos^2(2x)-\sin^2(2x)\] and now the first term is \[\cos^2(2x)=(\cos^2(x)-\sin^2(x))^2=\cos^4(x)-2\sin^2(x)\cos^2(x)+\sin^4(x)\]

OpenStudy (anonymous):

now lets see what happens with the \[\sin^2(2x)\]

OpenStudy (anonymous):

yeah it works out just as you want it to

OpenStudy (anonymous):

\[\sin(2x)=2\sin(x)\cos(x)\] and therefore \[\sin^2(2x)=(2\sin(x)\cos(x))^2=4\sin^2(x)\cos^2(x)\]

OpenStudy (anonymous):

subtract that off from the thing i wrote above and you get exactly what you want

OpenStudy (anonymous):

you are truly my savior, thank you! Now, do you think you will be able to also help me with the second problem?

OpenStudy (anonymous):

sure did you get the last one because i wrote it out in stages? it is clear right?

OpenStudy (anonymous):

Very clear. I got it.

OpenStudy (anonymous):

second one is \[\sec^2(x)=\frac{2\sec(x)}{\sec(x)+1}\]?

OpenStudy (anonymous):

\[\sec ^{2 }u/2 = 2secu / secu + 1\]

OpenStudy (anonymous):

Yes, you are right except for the left side should read sec squared times the quantity of X divided by 2.

OpenStudy (anonymous):

I'm sorry, don't mind me using U's instead of X's.

OpenStudy (anonymous):

oh it is \[\sec^2(\frac{x}{2})\] on the left?

OpenStudy (anonymous):

COrrect! Yes.

OpenStudy (anonymous):

oh lord. ok now i have to think. but before we think, lets write out what the right hand side really is \[\frac{\frac{2}{a}}{\frac{1}{a}+1}=\frac{2}{a+1}\] and so \[\frac{2\sec(x)}{\sec(x)+1}=\frac{2}{\cos(x)+1}\]

OpenStudy (anonymous):

Got it.

OpenStudy (anonymous):

good because that actually means we are done!

OpenStudy (anonymous):

That equals the left side?

OpenStudy (anonymous):

\[\cos(\frac{x}{2})=\sqrt{\frac{1+\cos(x)}{2}}\]

OpenStudy (anonymous):

so \[\sec(\frac{x}{2})=\sqrt{\frac{2}{1+\cos(x)}}\]

OpenStudy (anonymous):

and therefore \[\sec^2(\frac{x}{2})=\frac{2}{1+\cos(x)}\] just as we hoped

OpenStudy (anonymous):

you are awesome. This was so helpful. Thank you so, so much.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Is there a way I could find you if I needed help in the future? Luckily I only have one more month of this semester until Trig is finished.

OpenStudy (anonymous):

Thanks again.